### v2

# The Writing is on the Wall:

Abstract
Although rollups have attracted significant attention, there is limited empirical research on their performance under high load. To address this, we present a data-driven analysis of the transaction surge in late 2023 and early 2024, attributed to inscriptions—a method for recording data on the blockchain. Initially introduced on Bitcoin, inscriptions enable the representation of NFTs or ERC-20-like tokens without smart contracts, and have since expanded to other blockchains. This paper examines inscription-related transactions on Ethereum and major EVM-compatible rollups, assessing their impact on scalability during transaction surges. Our results show that, on certain days, inscriptions accounted nearly 90% of transactions on Arbitrum and ZKsync Era, while 53% on Ethereum, with 99% of these inscriptions involving meme coin minting. Furthermore, we show that ZKsync and Arbitrum saw lower median gas fees during these surges. ZKsync Era, a ZK-rollup, showed a greater fee reduction than the optimistic rollups studied—Arbitrum, Base, and Optimism.
1 Introduction
Since the introduction of Bitcoin [25] and later Ethereum [42], blockchains have seen increased adoption [5, 29]. This growth has led to the development of Decentralized Applicationss (DApps), including Decentralized Exchanges (DEXs) [3, 13, 19], lending protocols [28, 31], Non-Fungible Tokens (NFTs)for art [14], and other applications such as supply chain management [30, 32], and decentralized governance [3, 19, 22, 23].
In distributed systems, including blockchains, a fundamental trade-off has been observed; a blockchain can only optimize two out of the following three essential factors: security, decentralization, and scalability [41]; this is the “blockchain trilemma”, a term attributed to V. Buterin [41]. Traditionally, underlying blockchain solutions known as Layer-1 (L1)such as Bitcoin and Ethereum have prioritized decentralization and security at the expense of scalability. To address this challenge, Layer-2 (L2)solutions, built atop an L1, have been proposed. L2 solutions aim to improve scalability and efficiency while relying on the L1blockchain to heritage security. Consequently, a variety of Ethereum Virtual Machine (EVM)-compatible chains have emerged, focusing on scalability without compromising decentralization and security. This is facilitated by executing transactions off-chain and subsequently storing results or state transitions on an L1blockchain.
However, many of these L2scaling solutions were not yet fully stress-tested until late 2023, when a spike in transactions occurred: most L2 EVM-compatible chains experienced their all-year heights of transaction volumes in November and December 2023. This sudden transaction spike was caused by a boom in what is known as inscriptions [1]. Inscriptions originated from a new idea that surged in the Bitcoin blockchain — ordinals [40]. Ordinals were first introduced in March 2023 by an anonymous developer known as Domo [6]. In short, this new system allows users to assign unique identifiers to individual satoshis, the smallest unit of Bitcoin, enabling each satoshi to be tracked and traded separately, based on their ordinal number. This concept adds a new dimension to Bitcoin transactions by introducing the ability to treat each satoshi as a NFT, enabling new applications such as collectible assets, verifiable digital scarcity, and more complex smart contract interactions on the Bitcoin network. This led to a notable surge in the number of transactions in mid-2023 due to a Fear-Of-Missing-Out (FOMO)effect, which subsequently caused an increase in transaction fees for all users [39, 40].
Subsequently, the concept of ordinals was extended to other blockchains, such as Ethereum, Arbitrum, Base, Optimism, and ZKsync Era, thus leading to the new phenomenon of inscriptions. They enable users to embed arbitrary data directly onto these blockchains. This process involves attaching data, such as text, images, or code, to transactions or smart contracts, effectively “inscribing” this information permanently into the blockchain. Users quickly started to use inscriptions to mint and trade tokens, instead of using the already available smart contract functionality. Similarly to what happened with Bitcoin ordinals, inscriptions triggered a significant increase in Transaction-per-Second (TPS)on these chains, reaching yearly transaction highs in December 2023. During this time, inscription transactions represented over $80$% of all network transactions [1]. Due to their apparent popularity, various marketplaces dedicated to the trading of ordinals and inscriptions began to pop-up [8, 17, 48]. In particular, even the leading crypto-exchange, Binance, supports ordinals and inscriptions [9].
During the inscription boom at the break of 2023 and early 2024, Zero Knowledge (ZK)rollups — a type of L2blockchain — witnessed notably lower gas costs compared to Ethereum, increasing the interest of users, as the cost of participating in the inscription boom became considerably more affordable compared to Bitcoin and Ethereum. However, some rollups experienced downtime of approximately $78$ minutes [16, 36]. On March 13, 2024, Ethereum underwent the Dencun upgrade [2], which introduced blobs as temporary data storage. Blobs decreased gas costs for Ethereum rollups, but also raised concerns about the impact on inscription-based tokens ownership, since they rely on out-of-the-blockchain applications and therefore presumably assume data availability [27]. Note that after the 18-day period, blobs may disappear from the Ethereum blockchain, as they are designed to support state verification for rollups. Consequently, inscription-related data might only be available off-chain in the indexer of their platform creator if the rollup chooses not to store all their transaction data.
We summarize the contributions of this paper as follows.
Causes of transaction surge at the end of 2023.  Our analysis reveals that the surge in transaction volume were primarily driven by the minting of new inscription-based meme coins. These tokens were later traded for a fraction of the costs compared to inscriptions on the Bitcoin or Ethereum networks, with prices in the range of then cents to one US dollar. We found that the inscription standards on these chains are based on the BRC-20 standards with some modifications to facilitate trading [35, 40, 46] and listing on NFTmarketplaces [8, 17, 48]. Specifically, on December 17, 2023, for several hours, over 96% of all transactions on ZKsync Era involved the minting of ZRC-20 sync tokens. Other blockchains also exhibited a high percentage of inscription-related transactions. Arbitrum recorded the highest percentage, with over 89%, followed by Ethereum at over 53%, Base at over 37%, and Optimism at the lowest, with over 35%.
Impact on EVMblockchain performance and cost efficiency.  Given that certain L2 solutions experienced downtime, and a significant portion of transactions during late 2023 and early 2024 were inscriptions, we seek to evaluate the performance of these L2 scaling solutions during this stress-test. We found that the increased transaction activity resulted in a decrease in gas fees, positively affecting the users of the rollup chains. With more L2 transactions compressed into a single transaction batch, the median costs of L2 transactions decreased. Unlike Ethereum, where gas fees increased during periods of transaction surges, median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and ZKsync Era. Additionally, ZKsync Era, as a ZK rollup, demonstrated a stronger reduction in fees than Arbitrum, an optimistic rollup.
Speculative purchases.  We observed that across all chains, most on-chain inscription actions were for minting inscriptions rather than selling them. For example, the percentage of inscription minting activities were as follows: Arbitrum at 99.49%, Base at 99.52%, Ethereum at 98.71%, Optimism at 99.73%, and ZKsync Era at 99.6%. On ZKsync Era, the minting process was completed for three tokens: zrc-20 sync, era-20 bgnt and layer2-20 \$L2, which are traded on ZKsync Era for a fraction of the costs compared to the Ethereum and Bitcoin networks. Even with observed prices in the range of \$0.1 to \$1, the costs of minting and listing are covered for the seller. The trading activity for the era-20 bgnt token continued after the initial inscription boom.
Scientific reproducibility.  Gathering and processing blockchain data can be costly both financially and computationally, even for a single chain. Analyzing multiple chains, as in this study, increases complexity because most of the data are not even easily accessible. To promote scientific reproducibility and enable other researchers and companies to study blockchain data and inscriptions, we plan to make our datasets and scripts publicly available.
2 Background on Rollups
Scalability challenges and gas price fluctuations in blockchain networks are typically addressed through Layer-2 (L2)solution. While, L1is the foundational layer for blockchains, L2is a secondary layer that provides scalability and relies on L1for security. This also introduces rollups, a type of non-custodial L2blockchain, and provides the necessary background for this paper.
Layer-1 (L1).  This approach involves creating an entirely new blockchain with unique consensus mechanisms and dedicated physical infrastructure to maintain network consensus. Notable L1blockchains include Bitcoin, Ethereum, Solana, Avalanche, and Polkadot. L1blockchains typically prioritize decentralization and security, often at the expense of scalability. To address this limitation, L2-rollup solutions were proposed to help scale L1blockchains.
Layer-2 (L2).  In contrast, L2scaling adopts a different strategy. It involves executing computations outside the main (L1) blockchain and then recording the results or final state on the underlying chain. Major types of L2solutions include state (payment) channels, plasma, and rollups. Plasma and state channels focus on moving both data and computation off-chain. Rollups, on the other hand, move computation and state storage off-chain while retaining compressed data for each transaction on the underlying chain.
Rollups.  They function as a form of L2scaling by performing computations off-chain and storing the results on the L1chain. They operate like blockchains, generating blocks and then recording batches of these blocks on the L1chain. However, in the rollup environment, operators are not inherently trusted. This lack of trust means that operators could potentially behave maliciously by halting block production, generating invalid blocks, withholding data, or engaging in other adversarial activities [10]. To mitigate these risks, rollups typically implement robust security measures to ensure the integrity and security of the system. In that regard, ensuring the correctness of the state in rollups involves two main approaches: optimistic and Zero Knowledge Proof (ZKP). Some rollups, such as ZKsync Era, employ ZKP, which involve computing cryptographic proofs to validate the correctness of the computations. On the other hand, rollups like Arbitrum and Optimism adopt an optimistic approach, which assumes all statements are true unless someone in the network contests them.
Sequencers.  These entities are responsible for aggregating, ordering, and rolling up the transactions to the L1chain. By bundling transactions, rollups effectively achieve cost savings in terms of gas fees.
Verifiers.  Operating as smart contracts on the L1chain, verifiers play a crucial role in validating the transactions stored by the sequencer. They ensure the correctness of the transactions, thereby maintaining the integrity of the overall system [15, 34, 45].
3 Inscriptions and Ordinals
Inscriptions and ordinals involve the recording of arbitrary data on the blockchain. On EVM-compatible chains, users encode HEX data into the transaction input call data (refer to column JSON in Table 1), usually also setting the transaction’s from and to attributes with the same user addresses. This structure constitutes a self-transfer made from a user to their own address.
3.1 Operation Types
There are various protocol standards that define the structure inscriptions. BRC-20 was the first standard proposed for Bitcoin, and includes three main types of operations: deploy, mint, and transfer, each encoded in a single transaction [39]. The protocol standards on EVM-compatible chains extend this set of operations by adding, for example, claim, list, buy, sell, among others.
Table 1: Example of inscriptions data recorded on-chain: protocol name, operation, token name (Tick), total supply, and maximum amount of tokens minted each round (Limit)
Protocol Tick Operation Total Supply Limit JSON data
zrc-20 sync deploy $21\times 10^{6}$ $4$"p":"zrc-20","op":"deploy","tick":"sync","amt":"21000000","limit":"4"}
zrc-20 sync transfer $4$"p": "zrc-20", "op": "transfer", "tick": "sync", "amt": "4"}
zrc-20 sync mint $4$"p":"zrc-20","op":"mint","tick":"sync","amt":"4"}
zrc-20 zksi deploy $21\times 10^{8}$ $1\times 10^{4}$"p":"zrc-20","op":"deploy","tick":"zksi","max":"2100000000","limt":"10000"}
zrc-20 zkss deploy $21\times 10^{6}$ $1\times 10^{3}$"p":"zrc-20","op":"deploy","tick":"zkss","max":"21000000","lim":"1000"}
zrc-20 zkss sell $1000$"p":"zrc-20","op":"sell","orderMessage":[{"amt":10000,"nonce":"62...",
zrc-20 zkzk deploy $21\times 10^{6}$ $21\times 10^{6}$"p":"zks-20","op":"deploy","tick":"zkzk","max":"21000000","lim":"21000000"}
era-20 bgnt deploy $21\times 10^{6}$ $5$"p":"era-20","op":"deploy","tick":"bgnt","max":"21000000","lim": "5"}
era-20 bgnt deploy $21\times 10^{6}$ $5$"p":"era-20","op":"deploy","tick":"bgnt","max":"21000000","lim": "5"}
era-20 bgnt list $5$"p":"era-20","op":"list","tick":"bgnt","amt":"250","price":"1500000000000000"}
era-20 bgnt buy {"p":"era-20","op":"buy","tx":"0xda..."}
layer-2 \$L2 claim $1000$"p":"layer2-20","op":"claim","tick":"\$L2","amt":"1000"}
Deploy.  The deploy action specifies the protocol name, token tick, total supply, and the maximum amount of tokens a user can mint (or claim) per transaction. Table 1 shows examples of inscriptions’ data recorded on-chain. For example, to deploy an inscription, a transaction containing a deploy action should be recorded on-chain, marking the initiation of an inscription event. In the case of a zrc-20 sync inscription (see first row in Table 1), the protocol specifies a total supply of 21 million inscriptions and a limit of 4 tokens minted per transaction.
Mint.  After the transaction that deploys the inscription persists in the chain, users can issue a mint action to actually mint (or in this case claim ownership) of the tokens. To initiate this, users need to issue a transaction with an input call data encoded to HEX code, specifying the protocol and tick that jointly identify the inscription-based token. In the same transaction, users provide the number of tokens they want to claim (refer to column Limit in Table 1) that must not exceed the maximum limit specified in deploy operation for the given protocol-tick pair. The off-chain indexer is responsible for ensuring the integrity and user balances of inscription-based tokens.
Claim.  Similarly to mint, it allows users to mint and claim ownership of a new inscription token. This operation is used, for example, within the layer2-20 standard to mint a new inscription token, \$L2. The \$L2 token is a multi-L2inscription token that its inscription creator allows to be minted on multiple blockchains. However, there is just one deploy operation that defines the token’s maximum supply. The off-chain indexer maintained by the creator is responsible for ensuring the integrity of data in these circumstances.
Transfer.  Once users successfully mint the inscriptions, they can transfer ownership of their inscription tokens to another address. To do so, they typically issue a transaction to another address and add the data formatting standard of a transfer to the input call data. The receiving address will then possess ownership of these inscriptions. The inscription token is also identified by a protocol standard and a token name (tick). Tokens with the same names (ticks) can be deployed multiple times using various protocol standards.
List.  It enables users to list their inscriptions on marketplaces [8, 9, 17, 48]. Users specify the quantity of tokens and their price (in Ether) that they are willing to receive in exchange for their inscriptions from interested buyers.
Buy.  This operation is used to purchase inscription tokens from marketplaces. It references the transaction that listed the inscription on the marketplace. The buy operation specifies only protocol standards, as the token name (tick), amount, and price are declared in the list operation.
Sell.  It is used for selling inscriptions, this operation specifies the protocols and the order message containing seller and signer details, as well as the amount of tokens being sold and their prices.
3.2 Comparison with NFTs and ERC-20s
In this section, we discuss the distinctions between inscriptions and established Ethereum Request for Comments (ERC) standards such as ERC-20 [11] and ERC-721 [12] concerning tokens and Non-Fungible Token (NFT), respectively.
Comparison with NFTs.  In the Ethereum blockchain, NFTs [33, 38] are created through smart contracts. Each user receives a unique token ID, affirming their ownership of a specific asset. These assets, which can be JPEG files or CryptoPunk images [33], are stored off-chain on a server or on the InterPlanetary File System (IPFS) [18].
In contrast to NFTs, inscription-based tokens do not rely on any smart contract and, thus, do not allow upgrades. The link to the asset file is inscribed into the transaction data. In a blockchain with a maximal native token supply, such as Bitcoin, the amount of inscription is bound by the blockchain network limits. In contrast, NFTsbased on smart contracts are free from such limitations, theoretically allowing for unlimited minting. Also, each inscription is allocated a position in the blockchain, which creates the opportunity to derive the additional value of the inscription from the location within the block.
Comparison with ERC-20.  The other application of inscriptions is to create ERC-20-style tokens. As inscription-based tokens are not reliant on smart contracts, they are not susceptible to the risk of smart contract upgrades, and their maximal circulating supply is declared once in the deploy operation. The minting and transfer of each new inscription-based token can be tracked directly on the blockchain and cannot be altered.
However, due to their lack of smart contract support, inscription-based tokens offer limited functionality. They are primarily utilized for speculative purposes and fall into the category of meme-coins [39]. Based on the inscriptions technology, it is possible to mint a single NFT-style token or a group of tokens with a predefined token supply in circulation. Nevertheless, it is the off-chain indexer, managed by the token’s creator, responsible for ensuring the data integrity of minted tokens. Trading of inscription-based tokens predominantly occurs within inscriptions marketplaces since they are not compatible with the ERC-20 standard required by DEXs [44].
4 Data Collection
Table 2: Data set used to analyze inscription events on five blockchains.
Chain Start date End date # of Issuers # of Blocks # of Inscriptions
[…]
We analyze the inscriptions data recorded on Ethereum and its major EVM-compatible rollups: ZKsync Era, Arbitrum, Optimism, and Base. From these L2s, ZKsync Era is a ZKrollup, whereas other chains are optimistic rollups.
We then collected blockchain data from various sources. For ZKsync, we use data obtained from an archive node and a Web3-compatible API for ZKsync Era. For Arbitrum, Base, Ethereum, and Optimism, we use data sources from Nansen BigQuery [26]. These data sources enable us to collect comprehensive data containing all information recorded on the evaluated chains. This includes data about blocks, transactions, and events (or logs triggered during the execution of smart contracts) specifically related to inscriptions. Table 2 provides an overview of the specifics of our dataset.
Our focus is on blockchain data associated with inscriptions. While a significant portion of inscriptions involves the issuance of self-transfer transactions (i.e., where issuers initiate a transfer to their own address), we adopt a broader perspective. This involves considering instances where inscriptions are added to the chain, irrespective of whether through a self-transfer or not. This is important to capture the total fees users spent on inscriptions.
To identify these inscriptions, we specifically search for transactions with input call data starting with “0x646174613a”, representing “data:” in ASCII. This criterion allows us to recognize inscriptions within the blockchain. The blockchains with the most number of transaction inscriptions in our dataset were ZKsync Era and Arbitrum. In total, we found $17054466$ transactions containing inscriptions in our ZKsync Era data set issued by $481687$ addresses and added to $2809054$ blocks. Similarly, we identified $16309035$ transactions containing inscriptions in Arbitrum, issued by $118544$ addresses and added to $3575299$ blocks. The average number of inscriptions per block are $6.07$ and $4.56$ for ZKsync Era and Arbitrum, respectively. We provide similar statistics for Base, Ethereum, and Optimism chains in Table 2. This dataset forms the basis for our empirical analysis of inscriptions.
Table 3: Inscription transactions per issuers.
Protocol # of Issuers Mean Std. Median Min Max
Arbitrum $118544$ $137.57$ $676.76$ $6$ $1$ $38050$
Base $79573$ $25.39$ $154.91$ $3$ $1$ $19674$
Ethereum $245008$ $26.50$ $245.60$ $3$ $1$ $67713$
Optimism $49112$ $30.04$ $168.05$ $3$ $1$ $19612$
ZKsync Era $481687$ $35.62$ $168.05$ $3$ $1$ $40770$
The distribution of inscription transactions per address is highly skewed towards a few participants, as shown in Table 3. Most inscription issuers initiated at least three transactions related to inscriptions. Arbitrum exhibited the highest median, with at least six transactions per issuer. Table 3 also highlights that a small number of issuers were responsible for a significant fraction of the total inscriptions. For instance, a single issuer accounted for $67713$ transactions to claim their inscriptions, followed by ZKsync Era with $40770$ transactions.
5 Empirical Analysis
In this section, we present our empirical analysis of the inscriptions recorded on Ethereum and its major EVM-compatible rollups: Arbitrum, Base, Optimism, and ZKsync Era. We characterize the inscription protocols, operations, tokens, and their trading dynamics. Then, we investigate the impact of sudden surges in inscription transactions on the blockchains’ performance.
(a) Absolute number of inscriptions.
(b) Percentage of inscriptions in comparison to the overall number of transactions.
Figure 1: Daily distribution of inscription transactions in our data set: (a) absolute number of inscription transactions; and (b) percentage of inscription transactions compared to the total number of transactions.
5.1 Overall Transactions
[…]
5.2 Inscriptions Characterization
Figure 2: Percentage breakdown of the top 3 inscriptions attributes — protocol, operations, and tokens — for the chains analyzed in our study: Arbitrum, Base, Ethereum, Optimism and ZKsync Era.
Our analysis indicates that inscription tokens are uniquely unified by a protocol-tick pair. Therefore, we aggregated data using this pair to identify the top inscriptions tokens. Fig. 2 illustrates the breakdown of inscription transactions based on protocol, operation, and the leading token. In the following, we provide our observations for each of these categories.
Protocols.  Our initial observation reveals, in (Fig. 2 row Top Protocols), that each blockchain contains a particular leading inscription protocol, similar to BRC-20 on Bitcoin [46, 40], constituting between 54% and 93% of transactions. In addition, blockchain-specific standards such as FAIR-20 for Arbitrum, ERC-20 for Ethereum, and ZRC-20 for ZKsync Era appear the most. The LAYER2-20 protocol is present across all studied L2s, being the dominant inscription protocol on Base and Optimism, and the second largest on ZKsync Era.
Operations.  Recall from §3 that both claim and mint operation types work similarly. Thus, if evaluated together they account for approximately 99% of all inscription transactions in all chains we studied. From Fig. 2 row Top Operations, a small number of transactions are attributed to the list operation, used for listing inscriptions for sale in an inscription marketplace, and the deploy operation, which declares the token on the chain along with its token supply. The low percentage of transfer, sell, and buy transactions suggests that inscription tokens are not yet actively traded on EVMchains.
Tokens.  From our analysis (Fig. 2, row Top Tokens), we observe that various protocols deploy and mint tokens with the same tick symbols. To account for this, we identify each token using a protocol-tick pair. Our findings indicate that Arbitrum, Base, and Optimism each have a dominant token, comprising between 74% and 89% of inscription transactions on their respective platforms. In contrast, the token distribution on Ethereum is more diverse, with the ERC-20 NODES token contributing approximately 32% of the transactions, while all other tokens individually represent less than 3.14%. On ZKsync Era, the top four tokens account for over 92% of inscription transactions, also indicating token distribution concentration when compared to other chains. Notably, the token LAYER2-20 \$L2 is present across all rollups and is also the dominant token on both Base and Optimism.
5.3 Inscription Trading
We analyze the trading behavior of two inscription tokens within the ZKsync Era: zrc-20 sync (the predominant token) and era-20 bgnt (a smaller but noteworthy token due to its completed minting process). Our analysis, presented in Fig. 3, shows their respective trading activities. Notably, era-20 bgnt has a higher level of trading activity compared to zrc-20 sync. Specifically, we observed $477$ transfer transactions for zrc-20 sync, along with the corresponding token median prices as shown in Fig. 3(a), and $1148$ buy transactions for era-20 bgnt, with corresponding median prices depicted in Fig. 3(b)
(a) Transfer transactions of zrc-20 sync.
(b) Buy transactions of era-20 bgnt.
Figure 3: Price per token at every transaction on ZKsync Era considering zrc-20 sync and era-20 bgnt tokens.
Our analysis shows that the price of zrc-20 sync oscillated between \$0.10 and \$0.4, while era-20 bgnt ranged from \$0.10 to \$0.8. The trading of the era-20 bgnt token continued from its minting throughout the entire studied period. As the trades occur in batches of 4 or 5 tokens, they provide minimal compensation to the seller to cover the gas fees necessary to mint and list the inscriptions. However, trading inscriptions at such low prices is feasible due to the low gas fees on rollups, which ranged at ZKsync Era from \$0.05 to \$0.25 (see Fig. 4).
Table 4: Inscription operations cost in GWei for major token.
Chain Token Operation Total (in ETH) Mean Std. Median Min Max
Arbitrum fair-20 fair mint $0.77$ $74.10$ $305.72$ $60.87$ 00 $488849.89$
Arbitrum layer2-20 \$L2 claim $0.03$ $40.12$ $18.83$ $34.77$ $0.57$ $500.85$
Base layer2-20 \$L2 claim $0.0016$ $1.03$ $8.27$ $20.0$ $2199.20$
Ethereum erc-20 nodes mint $0.19$ $176.70$ $7015.97$ $99.32$ $0.01$ $1351757.78$
Optimism layer2-20 \$L2 claim $0.0009$ $0.78$ $3.04$ $0.11$ 00 $1617.84$
ZKsync Era zrc-20 sync mint $88.86$ $12684.59$ $3795.42$ $12748.95$ 00 $1097718.16$
ZKsync Era era-20 bgnt mint $11.31$ $7114.49$ $2525.91$ $6292.83$ $0.32$ $244836.15$
ZKsync Era era-20 bgnt list $0.015$ $4839.49$ $9528.53$ $272.57$ 00 $60807.7$
ZKsync Era layer2-20 \$L2 claim $0.017$ $29.63$ $6.90$ $28.38$ $0.42$ $2001.14$
Next, we analyze how spikes in inscription transactions influence gas fees across the network. Table 4 presents a breakdown of fees incurred during inscription operations. Recall that, in the first transaction, deploy, the maximum number of tokens that can be minted or claimed in a single transaction is specified. Then, we calculate the average fee per single token for each operation. We found that the mint operation is more efficient than claim, although both have a similar functionality. The interesting comparison provides claim operation for layer2-20 \$L2 token, as it is operating across all rollups. Its mining (claiming) operation had a lower median fee on Optimism and Base, followed by ZKsync Era and Arbitrum.
5.4 Impact on Gas Fees
(a) Transaction count.
(b) Median transaction cost.
Figure 4: Daily transaction count and median transaction cost in USD paid by the user: (a) transactions count; and (b) median transaction cost.
(a) L2 rollups.
(b) Ethereum mainnet (L1).
Figure 5: Distribution of gas fees in GWei relative to transaction count: (a) L2 rollups; and (b) Ethereum mainnet.
This section analyzes the impact of daily transaction volumes on the median gas fee per transaction. Unlike L1blockchains, rollups are expected to offer lower gas fees with increasing transaction counts due to their scaling mechanism. As more L2transactions can be compressed into a single batch stored in the L1 network, gas fees tend to decrease. In particular, spikes in daily transaction counts on specific days resulted in a decrease in median gas fees per transaction for other users, as shown in Fig. 4. Fees on L2rollups are linked to compression and are also affected by the fees of L1Ethereum.
Further analysis of the relationship between transaction count and median gas fee is presented in Fig. 5. Generally, gas fees with more transactions on rollups either remain unchanged (Base, Optimism) or decrease (Arbitrum, ZKsync). Conversely, on Ethereum, gas fees tend to increase with higher transaction counts. Additionally, ZKsync demonstrates higher efficiency in reducing gas fees compared to Arbitrum, attributable to its ZKrollup design, while Arbitrum, like Base and Optimism, is an optimistic rollup. Thus, unlike L1 blockchains, gas costs on ZKrollups decrease with a larger transaction volume, as observed on November 17, December 5, December 17, and December 20 on ZKsync Era.
6 Discussion
In this section, we discuss various aspects of inscriptions, starting with the impact of blobs on inscription ownership. We then explore whether inscriptions are beneficial or detrimental to the blockchain ecosystem. Finally, we address the potential for pump-and-dump schemes [21, 43] associated with inscriptions.
Impact of blobs on inscription ownership.  As inscriptions are appended to transactions as arbitrary data within the input call data, there is no assurance of perpetual availability of this data in the future. Blobs, initially designed to increase the efficiency of Ethereum rollups [2], were used to issue new inscription directly in Ethereum blockchains, called BlobScriptions [1]. Yet, blob data is intended to be stored by Ethereum nodes for 18 days, during which rollup operators or other parties verify their transactions. After this period, BlobScriptions disappear from the Ethereum blockchain and are only stored (and available) in the indexer of the creator protocol, outside the blockchain. Similarly, rollups are not required to store (outside of blobs) the transaction input data used by rollup-based inscriptions, raising questions about the ownership of inscriptions.
Inscriptions: good or bad?  A report by Binance on BRC-20 highlights that the unexpected surge in transactions resulted in an increase in transaction fees for many blockchains [35]. The report suggests that such an increase in fees can be seen as a natural progression in blockchain adoption. Therefore, attribution of this spike solely to BRC-20s is considered irrelevant. More concerning, however, is that other protocols, such as Arbitrum, experienced downtime of approximately 78 minutes [36], while ZKsync Era saw an increase in TPS, with nearly 96% of all transactions being related to inscriptions over the course of several hours. Thus, causing users to worry about delayed transaction inclusion. This sparked a broader debate within the blockchain community about whether ordinals and inscriptions truly benefit users.
Challenges and opportunities for trading inscription.  Inscriptions are not compatible with ERC-20 tokens. Therefore cannot be traded on traditional decentralized exchanges based on Automated Market Makers (AMMs)such as Uniswap [3], which is the base for trading tokens with low market capitalization. Today, inscriptions are listed at NFT-marketplaces in a limited-order book type, reducing their liquidity [8, 17, 48]. Also, for many inscription-based meme-coins, their minting process may still be ongoing, and new tokens could continue to be minted until the token limit specified in the deploy operation is reached. Furthermore, wrapping inscription-based meme coins into ERC-20 tokens could allow these inscriptions to be listed on DEXs, attracting meme-coin traders who are not familiar with NFTmarketplaces. We see this as the next evolution of inscriptions if they are intended to stay.
Hope or Hype?  During minting periods, platforms utilized various channels to engage users, including communities, influencers, and specific chains that amplified them. Consequently, numerous users issued inscriptions across different platforms [8, 17, 48]. Understanding user expectations, often influenced by FOMO, regarding inscriptions is crucial. For instance, are inscriptions more about hope or hype? This question relates to whether inscriptions are a novel concept or just a passing trend. In the case of Bitcoin, there was a significant surge in interest and user participation during the BRC-20 movement, followed by a subsequent decline in enthusiasm over the following months [39, 40]. Our work shows similar results for EVM-compatible chains: inscriptions peak initially but witness a decline in user interest over time, as shown in Section 5.1. The evolution of inscriptions fuels discussions on innovative features and integration with tokens and NFTs. Despite potential profitability, marketplace platforms encounter challenges in translating inscriptions into substantial profits for users, particularly across different chains. Due to the nature of inscriptions, which rely on off-chain applications, there is currently no mechanism to prevent users from purchasing previously sold inscriptions. Additionally, users can utilize software programs to automatically claim the majority of inscriptions during the minting or claiming period. Hence, should inscriptions be distributed directly to wallets, similar to traditional airdrop methods [24]?
Addressing these questions could potentially improve inscriptions, transforming them into a promising addition to the blockchain ecosystem.
7 Related Work
In this section, we review the existing literature and related works pertinent to our study on inscriptions within EVM-compatible chains.
Social phenomena on financial markets.  The impact of social phenomena on financial markets has received some research attention recently. In traditional finance, the GameStop short-squeeze event from 2021 highlighted how online social movements can lead to significant shifts in trading behavior and even impact established financial institutions [47]. Later, various authors analyzed this unprecedented event in more detail. Vasileiou et al. [37] found a causal link between the stock performance in early 2021 and the Google Trend Index for the stock, thus showing the impact of online interest in the stock in its price performance. Following this, Anand and Pathak [4] and Zheng et al. [47] focussed their analysis on the activity of the subreddit r/wallstreetbets (where the majority of the social activity related to GameStop was occurring during that time). Their findings further corroborate the link between stock performance and the subreddit’s social media activity and sentiment.
Social phenomena on decentralized finance.  Even though the introduction of ordinals and inscriptions is fairly recent, there has been some work towards understanding these phenomena. Wang et al. [39, 40] conducted a comprehensive analysis of BRC-20 tokens within the Bitcoin network. Their investigation revealed a notable surge in interest and user participation during the BRC-20 movement. The study involved a comparison between these BRCs and the widely-used Ethereum ERC-20 tokens, employing metrics such as average price return, volatility, and other relevant factors. The study highlighted a significant influx of users into the BRC market within a month, coupled with a subsequent decline in enthusiasm over the following months. In essence, users displayed a tendency to lose interest after the initial surge. In the broader context, the research emphasized that, despite the hype, BRC-based tokens constituted only a modest fraction of the overall market size when juxtaposed with ERC-like tokens on the Ethereum platform. Our paper also analyses the broader market dynamics of inscriptions during their boom, however, we expand the study to the top EVMchains instead of focussing on Bitcoin ordinals.
Li et al. [20] also looked at BRC-20 tokens within the Bitcoin network. However, they perform a comparative analysis between Bitcoin inscriptions and NFTs focussed on technical implementation details, function, and use cases. They additionally provide an in-depth historical overview of Bitcoin-related NFTs. While our paper also involves a characterization of inscriptions and their functions, our focus remains on EVM Layer 2s rather than Bitcoin.
Finally, Bertucci [7] analyzed the impact of ordinals on transaction fees in Bitcoin. They found that ordinal inscriptions typically have lower fee rates than regular transactions, which indicates careful fee-setting by ordinal users. They also concluded that the rise of ordinal transactions has increased non-inscription transaction fee rates and total block fees, thus increasing miners’ revenue. We do a similar fee investigation in our paper. Yet, we look at inscriptions in EVMchains instead of Bitcoin.
Novelty of our approach.  To the best of our knowledge, this is the first work to conduct an in-depth analysis of inscriptions in multi EVM-compatible chains and the user dynamics that arose from the inscription boom observed in the end of 2023 and early 2024. While Bitcoin originally set the stage for this user trend [39, 40], EVMchains have witnessed substantial adoption as well. The absence of analysis in this domain introduces a notable gap in the existing research landscape, which we address in this paper.
8 Conclusion
The inscriptions boom between December 2023 and January 2024 was the first real-life stress test for various EVM-compatible blockchains. Our study of Ethereum, Arbitrum, Base, Optimism, and ZKsync Era during this period revealed stark differences between L1, L2blockchains, and among rollups.
First, we observed that, on specific peak days, transactions related to inscriptions accounted for nearly 90% on Arbitrum and ZKsync Era, and over 53% on Ethereum, and almost 35% on Optimism and Base. The surge in on-chain inscriptions overwhelmed some rollups, causing them instability due to the increased transaction volume. This resulted in downtimes and delays in transaction processing.
We further found that 99% of these inscription transactions across all examined blockchains were driven by the claim or minting of inscription-based meme-coins, followed by a modest amount of trading activity, mostly associated with listing operations on marketplaces. All chains show a dominant token that represents more than 74% of all inscription transactions.
Looking at the impact of inscriptions on gas fees, we concluded that gas fees on rollups either remained stable or decreased during the transaction volume peak. Particularly, the median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and ZKsync Era. On the other hand, Ethereum experienced a rise in fees, which is not surprising. Unlike L1blockchains, rollups are designed to offer lower gas fees as transaction counts increase, due to their ability to compress multiple L2transactions into a single batch on the L1network.
Moreover, the future of inscriptions as a stable market for token trading in EVM-chains is still not cemented. For instance, Data Availability (DA)and User Interface (UI)issues can complicate interactions with inscriptions, highlighting how nascent the market still is. It is nevertheless an interesting trend and a lot of innovation can still occur to turn the space into a more mature one.
Finally, to promote scientific reproducibility and facilitate further research, we plan to make our datasets and scripts publicly available.
Acknowledgments
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.
Appendix 0.A Glossary
Following is a list of important notations used in this paper.
\@glotype@acronym@title